# BRIGHT: A globally distributed multimodal building damage assessment dataset with very-high-resolution for all-weather disaster response

### v1

@@ 4.2 Limitation of BRIGHT
However, the BRIGHT dataset also has some limitations. Firstly, our dataset involves both optical and SAR imagery covering the same locations. SAR images, in particular, can be distorted and stretched in certain areas. Despite thorough preprocessing, including manual alignment and cross-checking by multiple experts in EO data processing, minor alignment errors may still persist, as suggested by Table 3. Secondly, the building polygons in BRIGHT were manually annotated by expert annotators. Although manual labeling generally ensures high accuracy, minor errors in polygon boundaries are inevitable due to the complexity of building shapes and the variability in image resolution. These inaccuracies may slightly affect the performance of the models trained on BRIGHT. Then, the extent of damage to buildings is assessed by experts through visual interpretation of optical EO data. This process is susceptible to occasional misjudgments, contributing to label noise. Finally, although BRIGHT is globally distributed and richly geographically diverse, it has the problem of regional imbalance in the number of labels. Some of its events have more tiles and building numbers and thus are more dominant in training and evaluation, e.g., Turkey-Earthquake-2023 v.s. Hawaii-Wildfire-2023 in Table 2.
@@ 4.3 Significance of BRIGHT
Latency in responding to disaster events often arises from delays in interpretation methods and the acquisition of EO data (Ye et al., 2024). Traditionally, building damage proxy maps are generated through expert visual interpretation, a time-consuming process. Although ML- and DL-based methods offer automated and efficient approaches to building damage proxy mapping, they are often limited by the available training data. Large open-source optical datasets, though valuable, restrict models to optical imagery alone. The acquisition of clean optical EO data also requires sunlight and a cloud-free environment, which can result in delays of several days for certain disaster events such as typhoons and wildfires. As the first globally distributed multimodal dataset, BRIGHT encompasses both pre-event optical imagery and post-event SAR imagery. This unique combination overcomes the limitations of optical EO data by enabling models trained on BRIGHT to monitor disaster-stricken areas regardless of weather conditions or daylight. Compared to existing building damage datasets, BRIGHT offers several distinct characteristics: multimodal data, VHR imagery with resolutions better than 1 meter per pixel, coverage of six types of natural disasters, as well as man-made disasters, rich geographic diversity, and open access to the community. Due to these features, BRIGHT is anticipated to serve as a benchmark for many subsequent studies and practical disaster relief applications. BRIGHT facilitates research on building damage assessment. Looking forward, BRIGHT can be expanded to include additional data modalities such as height data, which would provide valuable topographic information to enhance disaster response strategies. Incorporating more bands in optical and SAR data could also broaden the applicability of the dataset.
@@
In addition to its primary application in building damage assessment, the BRIGHT dataset holds significant potential for other domains, particularly in multimodal EO data-related research. Firstly, BRIGHT can be directly used for as a building detection benchmark dataset with a rich geographic diversity and building patterns. Beyond that, researchers can also leverage BRIGHT to generate new datasets tailored to specific needs to explore new techniques including, but not limited to, image registration, land cover change detection, land cover mapping, height estimation, EO-driven visual question answering (VQA), improving the performance and reliability of models in complex environments. Moreover, BRIGHT can become an invaluable resource for the development of remote sensing foundation models (Wang et al., 2023; Hong et al., 2024). These emerging models, designed to generalize across various tasks and datasets, benefit immensely from the rich multimodal data and extensive geographic diversity offered by BRIGHT. By incorporating both optical and SAR data, BRIGHT enables the pre-training of more robust foundation models capable of handling the complexities of real-world remote sensing applications. This contribution is particularly important as the field moves towards creating versatile, scalable AI models that can be applied across different types of EO data and disaster scenarios (Li et al., 2024). We envision that BRIGHT, true to its name, will bring even a glimmer of bright to people in disaster-stricken areas by enabling more prompt and effective disaster response and relief.
@@
A prompt and comprehensive understanding of building damage following a disaster is critical for effective disaster response and relief operations. Currently, large-scale optical datasets support the development of automated ML/DL methods, but their use is limited by weather conditions and daylight availability. In contrast, all-weather disaster response can be facilitated by leveraging multimodal EO data. In this paper, we introduced BRIGHT, the first globally distributed, multimodal dataset with open access to the community, covering 12 natural and man-made disaster events. BRIGHT includes pre-event optical imagery and post-event SAR imagery with very high spatial resolutions (better than 1 meter). Alongside the dataset, we evaluated several state-of-the-art models on BRIGHT. The experimental results demonstrate the effectiveness of BRIGHT in mapping building damage across unseen events. These findings provide both performance baselines and valuable insights into model design for future research. BRIGHT is an ongoing project, and we remain committed to continuously enhancing its diversity and quality by incorporating new disaster events and refining the existing data. Our objective is to further improve BRIGHT’s utility for future research within the community and for practical disaster response applications at regional, national, and international levels.

### v2

@@ 5.1 Limitation of BRIGHT
However, the BRIGHT dataset also has some limitations. Firstly, our dataset involves both optical and SAR imagery covering the same locations. SAR images, in particular, can be distorted and stretched in certain areas. Despite thorough preprocessing, including manual alignment and cross-checking by multiple experts in EO data processing, minor alignment errors may still persist, as suggested by Table 4. Secondly, the building polygons in BRIGHT were manually annotated by expert annotators. Although manual labeling generally ensures high accuracy, minor errors in polygon boundaries are inevitable due to the complexity of building shapes and the variability in image resolution. These inaccuracies may slightly affect the performance of the models trained on BRIGHT. Then, the extent of damage to buildings is assessed by experts through visual interpretation of optical EO data. This process is susceptible to occasional misjudgments, contributing to label noise. Finally, although BRIGHT is globally distributed and richly geographically diverse, it has the problem of regional imbalance in the number of labels. Some of its events have more tiles and building numbers and thus are more dominant in training and evaluation, e.g., Turkey-Earthquake-2023 v.s. Hawaii-Wildfire-2023 in Table 2. Although we used an additional event-wise evaluation method, the effect of events, which account for a large percentage of the sample, on the model during the training phase is still not negligible. It may affect the generalizability of the trained model in real disaster scenarios.
@@ 5.2 Significance of BRIGHT
Latency in responding to disaster events often arises from delays in interpretation methods and the acquisition of EO data (Ye et al., 2024). Traditionally, building damage proxy maps are generated through expert visual interpretation, a time-consuming process. Although ML- and DL-based methods offer automated and efficient approaches to building damage proxy mapping, they are often limited by the available training data. Large open-source optical datasets, though valuable, restrict models to optical imagery alone. The acquisition of clean optical EO data also requires sunlight and a cloud-free environment, which can result in delays of several days for certain disaster events such as typhoons and wildfires. As the first globally distributed multimodal dataset, BRIGHT encompasses both pre-event optical imagery and post-event SAR imagery. This unique combination overcomes the limitations of optical EO data by enabling models trained on BRIGHT to monitor disaster-stricken areas regardless of weather conditions or daylight. Compared to existing building damage datasets, BRIGHT offers several distinct characteristics: multimodal data, VHR imagery with resolutions better than 1 meter per pixel, coverage of six types of natural disasters, as well as human-made disasters, rich geographic diversity, and open access to the community. Due to these features, BRIGHT is anticipated to serve as a benchmark for many subsequent studies and practical disaster relief applications. BRIGHT facilitates research on building damage assessment. Looking forward, BRIGHT can be expanded to include additional data modalities, such as height data, which would provide valuable topographic information to enhance disaster response strategies. Incorporating more bands in optical and SAR data could also broaden the applicability of the dataset.
@@
In addition to its primary application in building damage assessment, the BRIGHT dataset holds significant potential for other domains, particularly in multimodal EO data-related research. Firstly, BRIGHT can be directly used as a building detection benchmark dataset with a rich geographic diversity and building patterns. Beyond that, researchers can also leverage BRIGHT to generate new datasets tailored to specific needs to explore new techniques including, but not limited to, image registration, land cover change detection, land cover mapping, height estimation, EO-driven visual question answering (VQA), improving the performance and reliability of models in complex environments. Moreover, BRIGHT is a potential resource for the development of remote sensing foundation models (Wang et al., 2023; Hong et al., 2024). These emerging models, designed to generalize across various tasks and datasets, benefit immensely from the rich multimodal data and extensive geographic diversity offered by BRIGHT. By incorporating both optical and SAR data, BRIGHT enables the pre-training of more robust foundation models capable of handling the complexities of real-world remote sensing applications. This contribution is significant as the field moves toward creating versatile, scalable AI models that can be applied across different types of EO data and disaster scenarios (Li et al., 2024). We envision that BRIGHT, true to its name, will bring even a glimmer of bright to people in disaster-stricken areas by enabling more prompt and effective disaster response and relief.
@@
A prompt and comprehensive understanding of building damage following a disaster is critical for effective disaster response and relief operations. Currently, large-scale optical datasets support the development of automated ML/DL methods, but their use is limited by weather conditions and daylight availability. In contrast, all-weather disaster response can be facilitated by leveraging multimodal EO data. In this paper, we introduced BRIGHT, the first globally distributed, multimodal dataset with open access to the community, covering 14 natural and human-made disaster events. BRIGHT includes pre-event optical imagery and post-event SAR imagery with very high spatial resolutions (better than 1 meter). Alongside the dataset, we evaluated several state-of-the-art models on BRIGHT. The experimental results demonstrate the effectiveness of BRIGHT in mapping building damage across unseen events. These findings provide both performance baselines and valuable insights into model design for future research. BRIGHT is an ongoing project, and we remain committed to continuously enhancing its diversity and quality by incorporating new disaster events and refining the existing data. Our objective is to further improve BRIGHT’s utility for future research within the community and for practical disaster response applications at regional, national, and international levels.
@@
Appendix A Feature visualization
For the feature visualisation of buildings in different events in Figure 5c, we specifically employ DINOv2 (Oquab et al., 2024) to extract the high-dimensional features of buildings in different events. The corresponding reference maps are used to exclude background features. The features are then visualized using the t-SNE algorithm (van der Maaten and Hinton, 2008).
To visualize the spatial focus of models during building damage mapping, we apply Grad-CAM (Selvaraju et al., 2017), a gradient-based technique for generating class-specific activation maps. We perform Grad-CAM analysis on the ChangeOS model. Specifically, we compute the activation responses for the “Damaged” and “Destroyed” categories from the following components: the third or fourth ResNet block of the two branches of encoder, capturing mid- to high-level modality-specific features; the fusion module of the damage classification branch in the decoder, reflecting early-stage fusion and decision-making. These activations help reveal which spatial regions the model relies on when assessing different damage levels and how it utilizes information from optical and SAR modalities differently depending on the disaster context.
Appendix B Manual registration and estimating registration errors
Despite multiple rounds of meticulous registration and cross-validation of BRIGHT by several EO experts, registration errors between optical and SAR images could not be completely eliminated. Since we cannot directly obtain ground truth (i.e., actual ground control points), we adopt a proxy method for registration error estimation:
1.
Feature Point Selection: Select many feature points (e.g., 3000 points) from both the optical and SAR images using a keypoint detection algorithm.
- 2.
3.
Feature Matching: Perform feature matching between the extracted feature points across the optical and SAR images.
4.
Registration Error Computation: After completing the matching process, compute the pixel distances between the registered point pairs. These distances serve as a proxy for registration error.
5.
Outlier Removal: To improve the reliability of the estimation, we exclude matched points with a root mean square error (RMSE) greater than a certain threshold (e.g., 20 pixels). This threshold is set because after multiple rounds of expert corrections and cross-checking, large registration errors have already been eliminated. Points exceeding this threshold can be considered outliers.
| Event | RMSE (All) | RMSE (Building) |
| Beirut-EP-2020 | 1.365 | 1.365 |
| Bata-EP-2021 | 1.226 | 1.160 |
| Goma-VE-2021 | 0.966 | 0.975 |
| Les Cayes-EQ-2021 | 1.392 | 1.377 |
| La Palma-VE-2021 | 0.976 | 0.944 |
| Marshall-WF-2021 | 1.263 | 1.266 |
| Ukraine-AC-2022 | 0.899 | 0.882 |
| Turkey-EQ-2023 | 0.937 | 0.901 |
| Kyaukpyu-CC-2023 | 0.970 | 0.998 |
| Hawaii-WF-2023 | 1.124 | 1.124 |
| Morocco-EQ-2023 | 1.088 | 1.050 |
| Derna-FL-2023 | 1.267 | 1.255 |
| Acapulco-HC-2023 | 1.166 | 1.148 |
| Noto-EQ-2024 | 1.072 | 1.043 |
By following this method, we can estimate the registration errors between optical and SAR images in the dataset, providing a reliable foundation for further damage assessment analysis.
Figure of selected feature points (can be used as proxy ground truth for evaluating descriptors)
Appendix C Details of benchmark DL models
Here, we present some implementation details of our baseline model.
For UNet, the encoder has five convolutional blocks, each consisting of two 33 convolutional layers and two batch normalization layers. The number of channels in the five blocks are 64, 128, 256, 512, and 1024, respectively. The decoder contains four convolutional blocks, and the number of channels is set accordingly to the number of channels of the skip-connected features extracted by the encoder.
For DeepLabV3+, we applied ResNet-50 initialized with ImageNet pretraining weights (He et al., 2016) as the encoder. We then modified the input channel of the encoder so that the network can predict damage proxy maps directly from the optical and SAR stacked images.
For SiamAttnUNet, we followed the settings in its original literature (Adriano et al., 2021).
For SiamCRNN, we applied pseudo-siamese ResNet-18 as encoder and four ConvLSTM layers (Shi et al., 2015) with 33 convolutional kernel size and a hidden dimension of 128, as the decoder.
For ChangeOS, we used the ResNet-18 initialized with ImageNet pretraining weights as the encoder. We modified the encoder to a pseudo-siamese structure to extract features from pre-event and post-event images with different modalities.
For the ChangeMamba family of models (Chen et al., 2024), we chose one of them, MambaBDA-Tiny, for our experiments. We modified the encoder to a pseudo-siamese structure to extract features from pre-event and post-event images with different modalities.
Appendix D Details of post-processing methods
To further refine the raw predictions from deep learning models and enhance the quality of damage proxy maps, we apply three post-processing techniques: test-time augmentation (TTA), object-based majority voting, and model ensembling in Section 4.3. This section provides an overview of these methods and details the specific implementation used in our study.
Test-time augmentation is a widely used technique to improve model robustness by applying transformations to the input images at inference time. Instead of making a single prediction per image, multiple augmented versions of the same input are passed through the network, and the resulting predictions are aggregated. This reduces model sensitivity to spatial variations and increases prediction stability. In our study, we apply rotation (90°, 180°, 270°) and horizontal/vertical flipping to the input images during inference. The network produces a set of predictions for each augmented version, which are then aggregated at the logit level by summing the outputs before applying the softmax function.
Object-based post-processing aims to enforce spatial consistency by considering entire building instances rather than making independent pixel-wise predictions. Instead of classifying each pixel separately, the final label is determined based on a majority vote across the entire building object, leading to more coherent and reliable results. We follow the setup from (Zheng et al., 2021), where each building is treated as an independent object, and a weighted majority voting scheme is applied. The weighting is determined by the inverse of class proportions in the training set, meaning that underrepresented damage categories (e.g., destroyed buildings) have higher voting weights compared to dominant classes (e.g., intact buildings).
Model ensembling is a technique that combines predictions from multiple models to reduce uncertainty and improve generalization. By leveraging diverse models with different architectures, ensembling helps smooth out individual model biases and enhances the overall robustness of predictions. We perform ensembling by combining the results of three top-performing models: ChangeOS (Zheng et al., 2021), DamageFormer (Chen et al., 2022a), and ChangeMamba (Chen et al., 2024). Each model independently generates a damage proxy map, and the final prediction is obtained by averaging their logits before applying the softmax function.
Appendix E Details of unsupervised domain adaptation methods
| Events | AdaptSeg | AdvEnt | CLAN | PyCDA | FDA |
| Beirut-EP-2020 | 35.55 | 33.07 | 37.25 | 32.15 | 35.89 |
| Bata-EP-2021 | 37.54 | 35.20 | 36.98 | 30.01 | 36.31 |
| Goma-VE-2021 | 52.49 | 51.54 | 56.91 | 49.67 | 52.46 |
| Les Cayes-EQ-2021 | 37.11 | 36.86 | 39.47 | 33.40 | 33.90 |
| La Palma-VE-2021 | 30.54 | 30.04 | 30.10 | 32.76 | 30.50 |
| Marshall-WF-2021 | 37.20 | 35.01 | 34.35 | 30.14 | 33.93 |
| Ukraine-AC-2022 | 32.36 | 32.65 | 32.21 | 32.92 | 35.03 |
| Turkey-EQ-2023 | 33.69 | 34.54 | 36.68 | 34.58 | 36.82 |
| Kyaukpyu-CC-2023 | 34.62 | 29.22 | 30.83 | 31.31 | 36.71 |
| Hawaii-WF-2023 | 36.52 | 42.17 | 33.36 | 34.30 | 36.19 |
| Morocco-EQ-2023 | 37.93 | 37.42 | 39.04 | 30.55 | 38.76 |
| Derna-FL-2023 | 36.69 | 35.31 | 35.24 | 37.31 | 38.53 |
| Acapulco-HC-2023 | 27.05 | 27.32 | 27.73 | 30.11 | 31.68 |
| Noto-EQ-2024 | 35.49 | 35.65 | 33.99 | 25.44 | 31.39 |
| Average | 36.05 | 35.43 | 36.01 | 33.19 | 36.29 |
To evaluate the effectiveness of unsupervised domain adaptation (UDA) techniques in the cross-event zero-shot setting, we selected several representative methods from the computer vision literature originally designed for semantic segmentation tasks. The following methods are implemented:
AdaptSeg (Tsai et al., 2018): an adversarial training framework that aligns output space distributions between source and target domains.
AdvEnt (Vu et al., 2019): a variant of adversarial adaptation focusing on entropy minimization and structured output alignment.
CLAN (Luo et al., 2019): a category-level alignment method that selectively aligns class-specific features.
PyCDA (Lian et al., 2019): a curriculum domain adaptation framework that iteratively refines pseudo-labels across domains.
FDA (Yang and Soatto, 2020): a simple style-transfer-based approach that transfers source images into the target domain appearance via frequency mixing.
All methods are implemented using DeepLabV3+ as the backbone architecture. This choice is motivated by the fact that DeepLab-based networks are commonly used in the official implementations of these UDA methods, and many codebases are tightly coupled with DeepLab’s structure. Therefore, adopting DeepLabV3+ as the shared backbone ensures compatibility with existing implementations and avoids the need for extensive code modification. It also provides a fair and consistent basis for comparison across all methods.
In our experimental setup, each test disaster event is treated as the target domain, and only its unlabeled optical-SAR image pairs are used for adaptation. Importantly, no labeled target domain data is used for training or model selection, in line with our zero-shot assumption. This setup is fundamentally different from conventional UDA protocols in the computer vision community, where a target domain validation set is typically available to select the best model checkpoint, and the target test set is held out purely for evaluation. In contrast, our setting is designed to reflect real-world disaster response scenarios, where it is realistic to obtain unlabeled imagery, but no ground-truth annotations are available at adaptation time. Therefore, model selection must rely solely on source-domain feedback, which introduces additional challenges and more closely aligns with the operational constraints of emergency response.
The hyperparameters for each method closely follow the original papers to preserve method fidelity. Below are additional implementation details specific to certain methods:
FDA: For each training iteration, we generate frequency-transferred source images using FDA and mix them 1:1 with the original source images. The combined dataset is used to train the model from scratch.
PyCDA: Training proceeds in two stages. First, the model is trained on the source domain for 30,000 iterations. In the second stage, the adaptation module is applied to the target domain using pseudo-labeling and uncertainty-aware refinement as proposed in the original work.
For all methods, the same data preprocessing pipeline and training schedule (e.g., learning rate, batch size) are used unless otherwise specified in the original method, like those supervised models. The final evaluation is conducted on the target domain’s test set, and performance is reported in terms of mIoU.
Appendix F Details of semi-supervised learning methods
| Events | MT | CCT | GCT | CPS |
| Beirut-EP-2020 | 32.91 | 35.68 | 35.48 | 35.46 |
| Bata-EP-2021 | 37.94 | 36.61 | 38.25 | 36.96 |
| Goma-VE-2021 | 56.46 | 57.68 | 53.52 | 47.49 |
| Les Cayes-EQ-2021 | 40.20 | 39.57 | 38.83 | 39.41 |
| La Palma-VE-2021 | 33.52 | 33.03 | 36.90 | 33.29 |
| Marshall-WF-2021 | 39.73 | 40.12 | 38.74 | 32.94 |
| Ukraine-AC-2022 | 39.11 | 36.34 | 36.44 | 31.34 |
| Turkey-EQ-2023 | 38.96 | 39.72 | 36.50 | 37.15 |
| Kyaukpyu-CC-2023 | 36.72 | 35.84 | 33.56 | 30.22 |
| Hawaii-WF-2023 | 54.10 | 48.93 | 49.99 | 47.64 |
| Morocco-EQ-2023 | 37.18 | 40.54 | 41.10 | 37.54 |
| Derna-FL-2023 | 37.55 | 41.07 | 40.45 | 40.28 |
| Acapulco-HC-2023 | 32.30 | 30.16 | 33.36 | 31.51 |
| Noto-EQ-2024 | 43.40 | 37.61 | 37.74 | 37.65 |
| Average | 40.00 | 39.49 | 39.34 | 37.06 |
In the one-shot cross-event transfer setting, we evaluate several representative SSL methods for semantic segmentation. These methods were originally developed for natural image domains and are adapted here to the multimodal EO-based building damage assessment scenario.
Mean Teacher (MT) (Tarvainen and Valpola, 2017) maintains an exponential moving average (EMA) of model weights to generate stable predictions for unlabeled data, enforcing consistency between the student and teacher outputs.
CCT (Ouali et al., 2020) applies perturbations to the input data and enforces consistency between different views of the same image using multiple decoder branches.
GCT (Ke et al., 2020) further enhances this idea by introducing guidance from a pre-trained model to regularize learning.
CPS (Chen et al., 2021) trains two separate networks that generate pseudo-labels for each other, encouraging cross-supervised learning on unlabeled data.
All methods also use DeepLabV3+ as the backbone network. This decision is consistent with their original implementations and allows for seamless integration with official codebases without the need for architectural modifications. It also ensures a fair and consistent comparison across methods.
In the setup, 13 labeled disaster events serve as the training and validation sets, while a single labeled sample from the test (target) event is provided to simulate a realistic one-shot adaptation scenario. In addition, the remaining unlabeled samples from the test event are made available and used to facilitate SSL.
Each model is trained in two stages. In the first stage, the model is trained for 10,000 iterations using only standard supervised losses, including cross-entropy and Lovász-Softmax on the labeled data from the source domain and the single labeled sample from the target domain. This stage serves to initialize the model with a stable representation before semi-supervised training begins. In the second stage, each SSL method applies its respective semi-supervised objective using the unlabeled data from the target domain. These include consistency regularization (e.g., Mean Teacher), cross-view perturbation training (e.g., CCT and GCT), or dual-network pseudo supervision (e.g., CPS). Hyperparameters for each method follow the configurations suggested in their original publications.
Final performance is evaluated on the test set of the target domain using mIoU as the primary evaluation metric.
Appendix G Introduction of UMCD methods and a more practical evaluation protocol
We further evaluate several recent UMCD methods on the BRIGHT dataset. These methods are designed to detect binary changes from paired multimodal imagery without relying on manual labels.
IRG-McS (Sun et al., 2021) constructs superpixel-based graphs for each modality and iteratively matches their structural relationships to detect changes across heterogeneous image pairs, followed by Markov co-segmentation for final change map generation.
SR-GCAE (Chen et al., 2022b) learns robust graph-based representations of local and nonlocal structural relationships in multimodal images using a graph convolutional autoencoder, enabling effective change detection without supervision.
FD-MCD (Chen et al., 2023) proposes a Fourier domain framework that analyzes both local and nonlocal structural relationships using graph spectral convolution and adaptive fusion, enabling robust unsupervised multimodal change detection.
AOSG (Han et al., 2024) constructs an adaptively optimized structured graph to capture patch-level structural features in multimodal images, iteratively refining change intensity measures by fusing self-change and cross-domain structural differences for accurate unsupervised change detection.
AGSCC (Sun et al., 2024a) translates multimodal images via a structure cycle-consistent image regression framework that enforces similarity in structural graphs across domains, using adaptive graphs and multiple regularization terms to robustly detect changes without supervision.
AEKAN (Liu et al., 2025) utilizes a superpixel-based Siamese AutoEncoder built on Kolmogorov–Arnold Networks (KAN) to extract latent commonality features between multimodal images, using reconstruction and hierarchical consistency losses to detect changes in an unsupervised manner.
The training and evaluation protocols commonly adopted in the UMCD literature are often limited to individual image pairs (shown in Figure 12), using the same scene for representation learning, hyperparameter tuning, and evaluation. While this is understandable due to the lack of large-scale public datasets in this field, such a setup fails to reflect real-world use cases and often leads to overfitting and overestimated performance. To address this, we introduce a standardized and practical evaluation protocol using the BRIGHT dataset. Specifically, we use the validation set from the standard ML split in Section 2.3 as the training set for UMCD methods, including any hyperparameter or threshold tuning. These models are then evaluated on the combined training and test sets, which are strictly held out during adaptation. This avoids overlap between tuning and evaluation scenes and better simulates the real-world deployment setting, where a model is expected to generalize to new, unseen disasters.
For evaluation, it is important to note that most of these methods were originally developed for land cover change detection, where changes correspond to transitions between semantic categories such as vegetation, water, or built-up areas. Applying them directly to disaster damage detection poses certain challenges. In particular, “Damaged” buildings do not always exhibit strong spatial or spectral signals, making them hard to distinguish in an unsupervised setting. However, “Destroyed” buildings often result in a complete change of land cover appearance (e.g., collapsed structures, debris, or ground exposure), which aligns better with the assumptions of existing UMCD methods. Therefore, in our evaluation, we treat buildings labeled as “Destroyed” in BRIGHT as the positive (changed) class, and all other regions (including intact, damaged, and background) as the unchanged class. This allows a fair adaptation of these methods to a disaster-relevant setting while respecting their original design.